OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form and improper fractions are expressed as mixed numbers. A baker used 4 2/3 packages of raisins when he was making muffins. Each package contained 5 1/4 pounds of raisins. How many pounds of raisins did the baker use?

OpenStudy (anonymous):

4 packages, weighing 5 1/4 pound each is 21 pounds in total. 2/3 of a package weighing 5 1/4 pound is the same as: \[(\frac{ 2 }{ 3 })(\frac{ 21 }{ 4 })=7\] Add these two weights together: 21+7=28 pounds total